### v1

# Designing Simple Mechanisms

Designing Simple Mechanisms
Shengwu LiI thank Eric Chen for valuable comments.
( Preliminary draft solicited for the Journal of Economic Perspectives
May 2, 2024)
Which mechanisms are simple to play? When is it easy for participants to see that a mechanism is incentive-compatible? I will start by explaining how and why economists came to ask these questions. Then I will discuss three recent answers, that capture different aspects of what makes a mechanism simple.
Early work in mechanism design treated mechanisms as metaphors. For instance, Hurwicz (1973) describes a “resource allocation mechanism” as follows:
Simplifying to the utmost, we may imagine each agent having in front of him a console with one or more dials to set; the selection of dial settings by all agents determines uniquely the flow of goods and services (trade vector) between every pair of agents and also each agent’s production (input-output vector), his “trade with nature.”
Hurwicz’s mechanism is an abstract representation of richer real-world institutions. The dials stand for purchases that the agents could make, bids that they could place, bargaining strategies that they could adopt, and so on. Later work converged on a standard abstraction, in the form of the revelation principle: When studying what might happen in an arbitrary equilibrium of an arbitrary mechanism, it is without loss of generality to study what can be achieved by incentive-compatible revelation mechanisms. In such mechanisms, each agent is asked to directly report their private information, and it is in their best interest to report truthfully (Gibbard, 1973; Dasgupta et al., 1979; Myerson, 1979). A mechanism is Bayesian incentive-compatible if reporting truthfully is a Bayesian Nash equilibrium. A mechanism is dominant-strategy incentive compatible (‘strategy-proof’) if reporting truthfully is always a best response, regardless of the other players’ behavior.
On the metaphorical interpretation, it does not matter whether participants understand that the mechanism is incentive-compatible. After all, people do not interact with the mechanism per se; instead, the mechanism represents aspects of their everyday economic transactions. People may be proficient at arranging those transactions—making purchases in shops, haggling over contracts—and yet not recognize that the mechanism is incentive-compatible when confronted with its abstract form.
The limits of this interpretation became clear when economists began to design the rules for real-world marketplaces. This started in the 1990s, with auctions for radio spectrum broadcast licenses and the National Resident Matching Program (NRMP), a clearinghouse that matches American medical students to their first jobs (Milgrom, 2000; Roth, 2002). The practice soon found myriad applications; for instance, economists designed systems to coordinate electricity generation, to match students to public schools, to swap donated goods between food banks, and to sell loans from the Bank of England to financial institutions (Wilson, 2002; Abdulkadiroğlu et al., 2005; Prendergast, 2022; Klemperer, 2010).
The practice of market design led economists to focus on a more literal interpretation of mechanisms, namely that mechanism design involves “explicitly analyzing the consequences of trading rules that presumably are really common knowledge” (Wilson, 1987). On this view, the rules of the mechanism are not shorthand for a vague decentralized process; rather, they capture the actual rules of the institution as understood by the participants. For instance, the rules for a multi-round auction specify what bids are permitted, what each bidder learns about the bids from past rounds, when the auction will end, and how the winner and the payments will be determined.
When we regard the mechanism as capturing the rules of a formal process, such as an auction or a school choice system, it makes sense to ask whether participants understand that the mechanism is incentive-compatible. This question is particularly pressing for the purposes of market design, for several reasons. First, market designers often consider novel design proposals that are unfamiliar to the intended participants—or worse, may be deceptively familiar, leading participants to adopt heuristics that are unsuited to the new rules. Second, in applications such as school choice, people participate in the mechanism exactly once, and thus cannot learn from experience. Third, even if the designer asserts that the mechanism is incentive-compatible, participants may distrust that claim. And that distrust may be justified: Before the intervention of economists, the NRMP claimed that its algorithm made it incentive-compatible for medical students to report their true preferences, even though this was false (Williams, 1995). [^1] At the time, the NRMP ran the hospital-proposing deferred acceptance algorithm, which is not incentive-compatible for the applicants. Similarly, when Google started selling Internet advertising via the Generalized Second-Price auction, it incorrectly claimed that it was a dominant strategy for participants to bid their value, asserting that its “unique auction model uses Nobel Prize-winning economic theory to eliminate […] that feeling you’ve paid too much” (Edelman et al., 2007). (This presumably referred to the 1996 prize to William Vickrey and James Mirrlees, for work that we will shortly discuss.)
Participants in real-world mechanisms do not always recognize that the mechanism is incentive-compatible. The modern NRMP algorithm is (for all practical purposes) incentive-compatible [^2] The NRMP uses the applicant-proposing deferred acceptance algorithm, modified to account for couples. The standard deferred acceptance algorithm is strategy-proof for the proposing side (Dubins and Freedman, 1981), but accounting for couples slightly undermines this property. Computational experiments indicate that the probability that an applicant can profitably deviate under the NRMP algorithm is about $1$ in $10,000$ (Roth and Peranson, 1999)., and the NRMP prominently advises that participants should rank jobs in their preferred order. Nonetheless, 17% of graduating medical students claim, when surveyed, to have submitted a non-truthful rank-order list to the NRMP (Rees-Jones, 2018). Rees-Jones and Skowronek (2018) recruited medical students who recently participated in the NRMP, to be subjects in an incentived experiment with the same NRMP algorithm. They found that a similar proportion (23%) of subjects submit non-truthful rank-order lists in the experiment. As further evidence, in strategy-proof mechanisms that match students to degree programs, some students make unambiguous mistakes. They rank a program without a scholarship above the same program with a scholarship, even though the scholarship is worth thousands of euros (Hassidim et al., 2021; Shorrer and Sóvágó, 2023).
Consequently, it is often not enough that a mechanism is incentive-compatible; participants have to see for themselves that it is incentive-compatible; the mechanism has to be simple. Simplicity bypasses the need for participants to trust the designer. Simplicity eases the cognitive cost of participation, which is just as real as other costs. Simplicity can level the playing field, in the sense of Pathak and Sönmez (2008), preventing unfair outcomes caused by unequal strategic sophistication. Finally, simplicity allows the designer to rely more confidently on the predictions of classical game theory, because that analysis depends on participants correctly responding to incentives. [^3] Li and Dworczak (2024) make a formal critique of this last justification. If participants are strategically unsophisticated, then instead of using a simple mechanism, the designer might profitably adopt complex mechanisms that confuse the participants.
In the rest of this article, I explain three challenges that participants may need to overcome, in order to recognize incentive compatibility. These are:
  1. 1.
    Thinking contingently about unobserved opponent moves.
  2. 2.
    Planning for their own future moves.
  3. 3.
    Reasoning about other players’ beliefs.
In each of these cases, I will explain formal criteria that capture the difficulty in question, and discuss mechanisms that alleviate that difficulty.
The three challenges above are not exhaustive of the ways that mechanisms can be complex. This article is not a comprehensive survey, and it is shaped by my idiosyncratic tastes and the limits of my expertise. I have omitted excellent papers for the sake of brevity.
1 Thinking contingently about unobserved moves
In his seminal paper, Vickrey (1961) invented [^4] It is more accurate to say that Vickrey reinvented the second-price auction. Second-price auctions were used to sell postage stamps to collectors as early as 1893 (Lucking-Reiley, 2000). the second-price auction: All participants simultaneously place sealed bids, then the highest bidder wins and pays the second-highest bid. Vickrey argued that the second-price auction is desirable because it is simple:
Each bidder can confine his efforts and attention to an appraisal of the value the article would have in his own hands, at a considerable saving in mental strain and possibly in out-of-pocket expense.
The second-price auction is strategy-proof. One can see this by analogy. In an ascending auction, the price starts low and gradually rises. At each moment, each bidder decides whether to keep bidding or to quit irrevocably. When just one bidder is left, that bidder clinches the object at the current price. It is obvious that you should keep bidding if the price is below your value, and should quit if the price is above.
Vickrey pointed out that the second-price auction is “logically isomorphic to” the ascending auction. In the ascending auction, each bidder essentially chooses when to quit. The bidder with the highest quit-price wins, and pays the second-highest quit-price. Quit-prices in the ascending auction are equivalent to bids in the second-price auction, so it is a dominant strategy to place a bid equal to your value.
Notice that Vickrey’s argument runs the revelation principle ‘in reverse’. The second-price auction is an incentive-compatible revelation mechanism. We have reasoned about that mechanism by studying a dynamic mechanism—the ascending auction—and arguing that two mechanisms are equivalent. But do real participants treat them as equivalent?
For auctions in the wild, it can be hard to tell whether participants are bidding ‘truthfully’, because we usually do not know their value for the object. To overcome this obstacle, Kagel et al. (1987) had lab subjects bid in auctions for a ‘virtual prize’, which had a different value for each bidder. Each subject knew their own value, in dollars, for the virtual prize. The winning bidder received cash equal to their value minus the price they paid. This method enables the experimenter to observe whether subjects are bidding truthfully. [^5] Outside of revelation mechanisms, whether a strategy is ‘truthful’ is somewhat a matter of labeling. In such cases, I use ‘truthful’ to refer to the strategy that the designer intends for the participants to adopt.
Kagel et al. (1987) randomized whether subjects participated in ascending auctions or in second-price auctions. In theory, these two formats are equivalent; and it is a dominant strategy to choose a bid (or quit-price) equal to your value. [^6] Moreover, the money lottery induced by bidding your value first-order stochastically dominates the lottery from any other bid. Thus, bidding truthfully maximizes expected utility regardless of your risk preferences. Thus, we might expect that in both formats subjects will bid truthfully, and that the clearing price will always be equal to the second-highest value.
The theoretical prediction was largely borne out in ascending auctions: Subjects rapidly figured out that truthful bidding is optimal, and in 76% of auctions the clearing price was at the second-highest value (plus or minus one increment of the price clock).
By contrast, in second-price auctions, subjects did not behave as theory predicted. Subjects often submitted bids far from their values, the clearing price was at the second-highest value (plus or minus one increment) in no more than 20% of auctions. Subjects played the second-price auction 30 times in a row, with their values drawn randomly each time, but prices remained far from the theoretical prediction.
In summary, Kagel et al. (1987) found that lab subjects rapidly converge on truthful bidding in ascending auctions, but make large and persistent mistakes in second-price auctions. This result is well-replicated. [^7] For replications, see Kagel and Levin (1993), McCabe et al. (1990), Harstad (2000), Li (2017), and Breitmoser and Schweighofer-Kodritsch (2022). Ascending auctions and second-price auctions are not equivalent in practice.
What did our theory miss? What makes it easy for people to see that ascending auctions are strategy-proof? What makes it hard to see this for second-price auctions?
Suppose a participant wants to check directly that truthful bidding weakly dominates some deviating bid in a second-price auction; then they have to compare the payoffs for each profile of opponent bids. To bid optimally, they have to understand that if a truthful bid of \$40 would win at a price of \$30, then deviating to bid \$50 would also lead to a win at \$30. And if a truthful bid of \$40 would lose, then a bid of \$50 could only win at prices above \$40. If the bidder does not carefully keep track of each contingency, then they might be tempted to bid above their value in order to raise their chance of winning. The mistake is that this raises their chance of winning only when the price is so high that winning is undesirable.
By contrast, suppose that a participant is bidding in an ascending auction, and values the object at \$40. The price is now \$30. If the participant continues to bid truthfully, then they will either win at a price between \$30 and \$40, or they will lose and have payoff zero. Thus, every possible payoff from bidding truthfully is at least zero. On the other hand, if they deviate to quit right now, then their payoff is zero for sure. [^8] We have ignored the possibility of ties. Formally, we are modeling the ascending auction as a mechanism in which we cycle between the active bidders in some fixed order, asking each to raise their bid by a dollar or to quit irrevocably. When just one bidder remains, that bidder wins and pays their last bid. The worst-case payoff from truthful bidding is at least as good as the best-case payoff from deviating, so the participant need not keep track of different contingencies to see that truthful bidding is optimal.
A symmetric argument establishes that the participant should quit when the price exceeds their value. Suppose that the participant values the object at \$40 and the price rises to \$41. If they quit, then their payoff is zero for sure. If instead they keep bidding, then their payoff is no more than zero, because now they can only win at prices above \$41.
In hindsight, the distinction between second-price and ascending auctions looks trivial: To play optimally in a second-price auction, you need to compare strategies case-by-case [^9] Unless you notice, as Vickrey did, the isomorphism to the ascending auction. In a lab experiment, Breitmoser and Schweighofer-Kodritsch (2022) found that subjects are more likely to bid truthfully in second-price auctions when the mechanism is framed in a way that makes the isomorphism salient., whereas to play optimally in an ascending auction, you need not make that comparison.
We can generalize this idea to any extensive-form mechanism. Strategy $S_{n}$ obviously dominates strategy $S^{\prime}_{n}$ if at any information set where $S_{n}$ and $S^{\prime}_{n}$ choose different actions for the first time, the worst-case payoff under $S_{n}$ is at least as good as the best-case payoff under $S^{\prime}_{n}$. [^10] This results in a partial ranking of strategies, that can be understood as a weakening of a standard monotonicity axiom for choice under uncertainty (Echenique et al., 2022). (When the worst case from $S_{n}$ is preferred to the best case from $S^{\prime}_{n}$, a case-by-case comparison is not needed.) As we just saw, in ascending auctions, the truthful strategy obviously dominates any other strategy. A mechanism is obviously strategy-proof (OSP) if the truthful strategy is obviously dominant (Li, 2017).
Some strategy-proof mechanisms are not obviously strategy-proof. At the moment you submit your bid in a second-price auction, bidding your value can lead to a zero payoff, whereas bidding above your value can yield a positive payoff. Thus, truthful bidding is not an obviously dominant strategy.
Every obviously strategy-proof mechanism is strategy-proof. In fact, a stronger claim holds: Every obviously strategy-proof mechanism is weakly group strategy-proof, meaning that no coalition of players can jointly deviate and all strictly benefit. We will prove the contrapositive. Suppose the mechanism is not weakly group strategy-proof. Then there exists a strategy profile and a deviating coalition, such that the deviation is strictly profitable for every member. Along the resulting path-of-play, consider the first coalition member to deviate from the truthful strategy. That player strictly benefits, so at that information set, one possible outcome from the deviating strategy is strictly better than one possible outcome from the truthful strategy. It follows that the truthful strategy is not obviously dominant, which completes the proof.
Obvious strategy-proofness pins down the extensive form of the ascending auction. That is, suppose we restrict attention to auctions that are efficient (allocating the object to a bidder with maximal value) and in which only the winning bidder makes payments. In that class, an extensive-form mechanism is OSP if and only if it is an ascending auction (Li and Lo, 2023, p. 576–582). [^11] Alternatively, the ascending auction can be characterized as the only strategy-proof mechanism that is incentive-compatible for the auctioneer (Akbarpour and Li, 2020). Under weaker restrictions, one can characterize OSP mechanisms as ‘personal-clock auctions’, a larger class that includes ascending auctions, descending-price reverse auctions, and some hybrid formats (Li, 2017). Roughly, in a personal-clock auction each participant chooses between clinching a sure thing (e.g. leaving empty-handed) and a tentative alternative with a price that gets steadily worse for the participant (e.g. bidding to buy the object at a price that increases each round).
The 2017 Incentive Auction conducted by the US Federal Communications Commission (FCC) transacted US\$19.8 billion of radio spectrum rights. It included a reverse auction that paid TV stations to relinquish their over-the-air broadcast rights. This design problem involved computationally hard interference constraints, which made it infeasible to run a Vickrey-Clarke-Groves mechanism (Leyton-Brown et al., 2017). Instead, the proposed auction format used a suite of state-of-the-art feasibility-checking algorithms to determine the allocation and payments. Because of this complexity, it was especially important for participants to see that the proposed auction was incentive-compatible.
The reverse auction proceeds, roughly, as follows: The FCC makes an initial offer to each TV station. Each TV station can either quit or continue. If it quits, then it remains on the air, retaining its broadcast rights with zero net payment. If it continues, then either it sells its rights at the current price, or the FCC makes a lower offer and the process repeats. It is an obviously dominant strategy for each TV station to bid truthfully, quitting once the price drops below its value for remaining on the air. This reverse auction is another example of a personal-clock auction. It belongs to the class of deferred-acceptance clock auctions studied by Milgrom and Segal (2020), which have useful implications for privacy and budget-balance.
You can see that the FCC’s reverse auction is incentive-compatible, even though I have not described how the offers are determined or when the auction stops. [^12] These details depend on those complicated feasibility-checking algorithms. Most bidders lacked the computational horsepower to check whether the algorithms were running as intended. This is no accident. A mechanism is OSP if and only if truth-telling can be seen to be dominant from a certain kind of partial description. Such a description specifies the sequences of queries that one participant might receive, the answers that are allowed, and how queries and answers map to possible outcomes (Li, 2017, Theorem 1). Notice that this description omits how the outcome depends on unobserved opponent moves; in this sense, OSP mechanisms formalize the idea of dominance that can be recognized without contingent reasoning.
Psychologically, it is harder to account for hypothetical contingencies, than to reason about observed events (Esponda and Vespa, 2014). Dynamic mechanisms can help participants to avoid mistakes, by resolving uncertainty about what other players have done. This connects to a larger literature on contingent reasoning (Martínez-Marquina et al., 2019; Cohen and Li, 2022; Esponda and Vespa, 2023), reviewed by Niederle and Vespa (2023).
The assumptions we make about preferences affect the structure of OSP mechanisms. Recent work has explored beyond auctions, characterizing OSP mechanisms for settings such as two-sided matching, social choice, and object allocation without transfers (Arribillaga et al., 2020, 2023; Ashlagi and Gonczarowski, 2018; Bade, 2019; Bade and Gonczarowski, 2017; Mandal and Roy, 2022; Thomas, 2020; Troyan, 2019). Other work has built tools to study OSP mechanisms, providing a ‘revelation principle’ (Mackenzie, 2020) and algorithms to construct them (Golowich and Li, 2021).
2 Planning for future moves
Suppose that we have a set of indivisible goods and a set of agents, and we want to allocate one good to each agent, without money transfers. One natural solution is to order the agents randomly, give the first agent their favorite good, then give the second agent their favorite good among those that remain, and so on iteratively. This algorithm results in an ex post efficient allocation (Bogomolnaia and Moulin, 2001).
We can ask agents to simultaneously submit rank-order lists of the goods, and then process those lists according to the above algorithm. The resulting mechanism—the static random priority mechanism (static RP)—is strategy-proof. However, it is not obviously strategy-proof. If you rank goods truthfully, you might receive your third-favorite, whereas if you deviate to rank your second-favorite good at the top, then you might receive your second-favorite good, which you strictly prefer.
Alternatively, we could approach agents one-by-one in random order, asking each to pick their favorite good among those that remain. The resulting dynamic random priority mechanism (dynamic RP) is obviously strategy-proof.
In practice, dynamic RP results in higher rates of truthful play than its static equivalent. In an incentivized laboratory experiment with four players and four goods, dynamic RP resulted in the dominant-strategy outcome in 93% of games, whereas static RP resulted in the dominant-strategy outcome in only 64% of games (Li, 2017, p. 3282). [^13] Dreyfuss et al. (2022) and Meisner and von Wangenheim (2023) argue that this disparity might be due to loss aversion rather than strategic mistakes.
Ascending auctions and dynamic RP are both intuitively simple. But some games with obviously dominant strategies are not intuitively simple. For example, take any chess position such that White can force a win, and consider the subgame that starts from that position. [^14] It is not known whether White can force a win from the opening position. World Chess Champion Bobby Fischer opined, “I think it’s almost definite that the game is a draw theoretically”. If White plays the win-forcing strategy, then by definition the worst-case outcome is that White wins. If White ever deviates from that strategy, the best-case outcome is that White wins. Thus, the win-forcing strategy is obviously dominant in the subgame. However, real chess players often fail to play the win-forcing strategy, even from positions that appear in an endgame tablebase (Anderson et al., 2017).
What makes dynamic RP simple, and chess complicated? Intuitively, to identify a win-forcing strategy in chess, one has to plan far in advance, consider different future contingencies, and then backward induct in order to judge the merits of each current move. By contrast, when playing dynamic RP, each player chooses exactly once, so forward planning is not required. Conventional game theory elides this distinction, because a strategy is a complete contingent plan, specifying not only what one does right now, but also what one will do at all future contingencies.
Pycia and Troyan (2023) formalized the idea of limited forward planning, as follows: We consider games of perfect recall, that is, the information sets are such that each participant ‘remembers’ all the past information sets that they encountered and all the past moves they took. Let $\mathcal{I}_{n}$ denote the family of information sets assigned to player $n$. Suppose that at each information set $I\in\mathcal{I}_{n}$, player $n$ can only foresee some of the other information sets. We use $F(I)\subseteq\mathcal{I}_{n}$ to denote the family of information sets foreseeable at $I$, and assume that $I\in F(I)$. A partial strategic plan for information set $I$, denoted $S_{n}^{I}$, chooses actions at all information sets in $F(I)$. The plan $S_{n}^{I}$ is $F$-dominant if, upon reaching information set $I$, the worst-case outcome from following $S_{n}^{I}$ is at least as good as the best-case from deviating to another plan that chooses differently at $I$. Specifically, the worst-case and the best-case are with respect to the actions of other players and also the actions that player $n$ takes at unforeseeable information sets, those in $\mathcal{I}_{n}\setminus F(I)$.
This approach results in a variety of incentive criteria, that depend on how we specify the foreseeable information sets $F(I)$. At one extreme, all information sets are foreseeable, that is, $F(I)=\mathcal{I}_{n}$, and $F$-dominance is equivalent to obvious dominance. On the other extreme, only the present information set is foreseeable, so $F(I)=\{I\}$, and we could call the resulting criterion strong obvious dominance. For example, in dynamic RP, picking your favorite object is strongly obviously dominant. Each player is called to play exactly once, so it does not matter that future information sets are unforeseeable.
Sometimes optimal behavior requires an intermediate degree of foresight. Consider a bidder who values the object at \$40. They are participating in an ascending auction, with the price rising by \$1 in each step. The price starts at \$1, and the bidder reasons, “If I quit now my payoff is \$0 for sure. But if I agree to bid \$1 and plan to quit at \$2, then my payoff could be \$$\$40-\$1=\$39$, and is at least \$0. So I’ll keep going for now.” Then the price rises to \$2, and the bidder thinks, “If I quit now my payoff is \$0 for sure. But if I keep bidding, planning to quit at \$3, then my payoff could be \$$\$40-\$2=\$38$, and is at least \$0.” Thus, by looking one step ahead at each point, the bidder is led to a pattern of behavior that reproduces the optimal strategy. Notice that the bidder is making a collection of partial strategic plans, and these plans are not consistent with each other. When the price is \$1, they plan to quit at \$2. But when the price reaches \$2, they revise that plan.
To formalize this idea, let the foreseeable information sets $F(I)$ consist of $I$ and every information set that is ‘one step ahead’ of $I$, and let us call the resulting incentive criterion one-step dominance. [^15] That is, where the same player is called to play, without any other such information sets in between. Observe that the ascending auction is one-step simple, meaning that truthful bidding can be induced by a collection of partial strategic plans (one for each information set), each of which is one-step dominant. Just a little bit of foresight is enough to play optimally in an ascending auction.
Let us return to the setting with indivisible goods and no transfers. As we saw, dynamic RP is OSP. But there are other OSP mechanisms in this setting; Pycia and Troyan (2023) characterized those mechanisms, and some of them are not intuitively simple, instead requiring participants to plan far into the future in order to see that the mechanism is incentive-compatible. They showed that strengthening the requirement from OSP to one-step simplicity rules out the counter-intuitive mechanisms.
It can take detailed contingent thinking to see that a revelation mechanism, such as the second-price auction, is incentive-compatible. Dynamic mechanisms, such as the ascending auction, can mitigate this difficulty by paring down the contingencies that participants need to consider. [^16] Glazer and Rubinstein (1996) formalize another sense in which dynamic games can be easier to understand than their static counterparts. But dynamic mechanisms can raise new difficulties, requiring participants to plan far in advance in order to understand that a given strategy is optimal. As we saw, the concept of partial strategic plans introduced by Pycia and Troyan (2023) provides a way to study this kind of complexity.
3 Reasoning about other players’ beliefs
The study of simplicity is not only about refining strategy-proofness. Even mechanisms without dominant strategies can vary in how simple they are.
In some mechanisms, each player has to reason about the other players’ beliefs. Consider bilateral trade with transfers; there is a seller who can produce an indivisible object, and a prospective buyer. The seller has a privately-known cost $C$ and the buyer has a privately-known value $V$. If a sale is made at some price $p$, then the buyer’s payoff is $V-p$ and the seller’s payoff is $p-C$. Otherwise both payoffs are 00.
One natural mechanism for bilateral trade is the double auction, studied by Chatterjee and Samuelson (1983): The seller and buyer simultaneously submit offers, $s$ and $b$ respectively. They transact if and only if $s\leq b$, at price $\alpha s+(1-\alpha)b$, for parameter $\alpha\in[0,1]$. Suppose for the moment that $\alpha=0.5$, so the mechanism splits the difference between the two offers. Then neither side has a dominant strategy. The seller’s utility-maximizing offer depends on the buyer’s offer $b$; if the seller knew $b$, then they would set $s=b$ if $C\leq b$, and set $s>b$ otherwise.
Suppose it is common knowledge that both players are rational, that is, they both make offers that maximize expected utility. Since both offers are made at the same time, the seller does not know the buyer’s offer $b$. Instead, the seller has to form beliefs about the distribution of $b$. Since the seller knows that the buyer is rational, the seller knows that the buyer’s choice of $b$ depends on the buyer’s value $V$ and on the buyer’s beliefs about the seller’s offer $s$. And that, in turn, depends on the buyer’s beliefs about the seller’s cost $C$, and the buyer’s beliefs about the seller’s beliefs about $b$. Thus, the seller’s offer depends on a second-order belief; what the seller believes that the buyer believes about $C$, as well as a third-order belief; what the seller believes that the buyer believes that the seller believes about $V$, and so on.
This reasoning process seems fantastical, and in practice will be beyond the capacity of many participants. [^17] In many theoretical analyses, we abstract away from this subtlety by studying Bayesian Nash equilibria and making strong assumptions about higher-order beliefs; for instance, that both buyer and seller have a common prior about the distribution of values. These simplifying assumptions are often useful. There is a recent literature that seeks to relax them (Bergemann and Morris, 2005; Bergemann et al., 2013). If participants are inexperienced, it seems unlikely that they will find the equilibrium of the double auction with parameter $\alpha=0.5$ from first principles.
Some mechanisms do not require participants to reason about higher-order beliefs. Suppose that we take the double auction, and instead set the parameter $\alpha$ equal to $1$. Then the seller is effectively making a take-it-or-leave-it offer of $s$ to the buyer, and the transaction occurs at a price of $s$ if and only if $s\leq b$. It is a dominant strategy for the buyer to set $b=V$, so the buyer does not need to think about the seller’s strategy. If the seller knows that the buyer will set $b=V$, then the seller’s problem reduces to choosing an offer $s$ to maximize $(s-C)P(s\leq V\mid C)$. Observe that the solution depends on the seller’s cost $C$ and on the seller’s belief about the buyer’s value $V$, but not on any higher-order beliefs. The seller can calculate their optimal offer using only their first-order beliefs, although they do not have a dominant strategy.
Börgers and Li (2019) proposed a new incentive criterion, picking out mechanisms in which participants can ‘play well’ using just their first-order beliefs. The proposal is as follows: Let us fix, for each player $n$, a set of possible utility functions $\mathcal{U}_{n}$ (e.g. the different values that the buyer might have for the object). A strategy for player $n$ is a function from $\mathcal{U}_{n}$ to distributions over actions (e.g. in the double auction, the actions are the feasible offers). A strategy is undominated if there is no other strategy that is always at least as good, and sometimes strictly better. A first-order belief $\mu_{n}$ for player $n$ is a distribution over the other players’ utility functions, that is $\mu_{n}\in\Delta(\mathcal{U}_{-n})$. For instance, the seller might believe that the buyer’s value is uniformly distributed on $\{0,1,2,\ldots,10\}$. A strategy $S_{n}$ is robust with respect to first-order belief $\mu_{n}$ if, assuming the other players’ utilities are distributed according to $\mu_{n}$, the strategy $S_{n}$ is a best response to every profile of undominated strategies for the other players. A mechanism is strategically simple if for every player $n$ and every first-order belief $\mu_{n}\in\Delta(\mathcal{U}_{-n})$, there exists a robust strategy with respect to $\mu_{n}$.
Every strategy-proof mechanism is strategically simple, but not vice versa. For example, the double auction with $\alpha=1$ is strategically simple. For the buyer, offering $b=V$ is a dominant strategy, and hence a robust strategy. Moreover, offering $b=V$ is the buyer’s only undominated strategy, so the seller has a robust strategy that sets $s$ to maximize $(s-C)P(s\leq V\mid C)$, which depends only on the seller’s cost $C$ and on the seller’s beliefs about the distribution of $V$, i.e. the seller’s first-order belief. By contrast, the double auction with $\alpha=0.5$ is not strategically simple. The buyer has many undominated strategies, and for non-trivial seller beliefs there is no robust seller strategy.
Börgers and Li (2019) proved that in the bilateral trade setting, all the strategically simple mechanisms have a special structure: The ‘leading’ player either chooses an offer $p$ from a set of permitted prices, or declines trade. If an offer of $p$ is made, then the ‘following’ player either accepts (and trade occurs at price $p$) or declines, in which case no trade occurs. This class includes the double auction with $\alpha=0$ and with $\alpha=1$, as well as variations that restrict the set of permitted prices.
More generally, Börgers and Li (2019) found that, under a richness condition, the strategically simple mechanisms can be characterized as ‘local dictatorships’. Roughly, this means that if we fix a profile of utility functions, and consider the restricted mechanism that includes only actions that are undominated at that profile, then in the restricted mechanism only one of the player’s actions determines the outcome.
The class of local dictatorships is subtle, and includes mechanisms that are not ‘dictatorial’ in the colloquial sense. One example of a local dictatorship is a voting mechanism in which each committee member submits a list ranking all alternatives and (at the same time) the chairperson selects two alternatives for a run-off vote. From the selected pair, the final outcome is the alternative that is ranked higher by a majority of the committee. (Suppose that all players have strict preferences over the alternatives.) Fixing a profile of utility functions, each committee member has only a single undominated action, namely, to submit their truthful rank-order list. Thus, in the restricted mechanism only the chairperson’s action determines the outcome.
4 Conclusion
Mechanisms can be complicated in many ways; challenging players to think contingently about other players’ moves, to plan for their own future moves, and to reason about other players’ beliefs. Thus, there are multiple ways to define what makes a ‘simple mechanism’. The right criterion depends on context; participants in designed mechanisms include schoolchildren, doctors, lumberjacks, fishermen, and telecommunications firms advised by game theorists (Athey and Levin, 2001; Marszalec et al., 2020; Bulow et al., 2009).
Why have formal simplicity criteria at all? After all, we know it when we see it. If we use our intuition to judge formal criteria, why not use intuition instead of formal criteria? I see two considerations that weigh against the intuitive approach. First, simplicity is not the only goal—there are often others, such as efficiency, fairness, or revenue. Unless we put simplicity on an equal mathematical footing, we cannot study these trade-offs systematically. Without formal simplicity criteria, we may fixate on other desiderata that are well-formalized. Second, our intuitive judgements about simplicity are distorted by economic training. Experts in mechanism design suffer from the curse of knowledge; it is hard to adopt the perspective of people who do not know what we know (Camerer et al., 1989). In order to produce scientific knowledge about simplicity, we must construct precise theories that can be tested with data.
There is a pressing need for experiments, both to test current theories about simplicity and to discover empirical regularities that future theories might explain. In testing theories, bear in mind that some tasks might be easy if done in isolation, but difficult when they are part of a broader strategic context. [^18] Moreover, in real-world mechanisms, participants must figure out their own preferences; each must devote “efforts and attention to an appraisal of the value the article would have in his own hands” (Vickrey, 1961), a complication that is often removed in laboratory experiments. Kagel and Levin (2016) survey experiments on auctions and Hakimov and Kübler (2021) survey experiments on matching mechanisms.
Much remains to be done. The criteria we have studied so far are not exhaustive; there are other dimensions on which mechanisms can be simple or complex. There could be better ways to formalize the same dimensions, that are more tractable, more closely track human behavior, or have firmer cognitive foundations. Moreover, simplicity comes in degrees, and it would be useful to compare the relative simplicity of different mechanisms, or to find ways to say that one mechanism is ‘as simple as possible’ given other constraints. Nagel and Saitto (2023) have recently made progress in this direction. Finally, another important direction is to study how to describe and explain mechanisms, to help participants see for themselves that the mechanism is incentive-compatible (Gonczarowski et al., 2023).